LASER INDUCED FLUORESCENCE SPECTROSCOPY OF RHODIUM FLUORIDE

\underline{ROMEY F. HEUFF}, WALTER J. BALFOUR, RUNHUA LI and ROY H. JENSEN, Department of Chemistry, University of Victoria, Victoria, British Columbia, Canada, V8W 3P6; ALLAN G. ADAM and SCOTT A. SHEPARD, Department of Chemistry, University of New Brunswick, Fredericton, New Brunswick, Canada E3B 6E2.

Rhodium monofluoride has been observed and spectroscopically characterized for the first time. RhF molecules were produced under jet-cooled conditions in a laser vaporization molecular beam source by the reaction of a laser-vaporized rhodium plasma with SF_{6} doped in helium, and studied with laser-induced fluorescence spectroscopy under both mediumand high-resolution. More than 25 LIF bands have been observed between 18 500 and 24 500 cm^{−1}and five of these have been recorded at 200 MHz resolution. All bands of appreciable intensity have been rotationally analyzed. The ground electronic level has Ω = 2, which is attributed to an inverted ^{3}Π state from the 2δ^{4}12σ^{1}6π^{3} electron configuration. The ground level rotational constants are B = 0.27245 cm^{−1}, D = 1.035x10^{−7} cm^{−1}. Very small ground level Λ - doublings are evident in the spectrum. Excited states having Ω = 1, 2 and 3 have been identified. Dispersed fluorescence spectroscopy from eleven excited levels has been used to locate a large number of low-lying vibronic states within the energy range up to 8000 cm^{−1}. A ground state vibrational interval of ~575 cm^{−1}is suggested.